OpenStudy (anonymous):

considering two numbers. the sum of two numbers is 26. one-fourth the larger number is equal to three times the smaller number. what are these two numbers?

OpenStudy (anonymous):

let the numbers be x and y and x>y x+y=26 x/4=3y x=12 y 12y+y=26 y=2 find x

OpenStudy (anonymous):

x=24. Thank you, I messed up the second equation so I was confused.